OpenStudy (anonymous):

16x^2+25y^2-128x+200y+256

OpenStudy (anonymous):

ellipse

OpenStudy (anonymous):

yes there is a formula that u have to use. Do u have it by any chance?

OpenStudy (anonymous):

ok so basically we have to change it into this form (x - h)²/a² + (y - k)²/b² = 1

OpenStudy (anonymous):

ok so what u do is lets gather the x and y terms 16x^2-128x+25y^2+200y+256

OpenStudy (anonymous):

ok so now we have to complete the square for the x's and the y's and we move over the 256 to the other side and remember whatvr u do to one side u gotta do to the other 16(x^2-8x+16)+25(y^2+8x+16)=-256+16*16+25*16

OpenStudy (anonymous):

Did u follow so far?

OpenStudy (anonymous):

hmm u disappeared on me